Hello all.
May you please help me answer this word problem? Thanks:
The height of a rectangle is half as long as its base.If its height is increased by 5 and its base is decreased by 2, its perimeter is equal to 60. Find the height and base of the rectangle.

Answer by chana(69)

The height of a rectangle is half as long as its base.If its height is increased by 5 and its base is decreased by 2, its perimeter is equal to 60. Find the height and base of the rectangle.
sol:
let height = b
base = l
as given :
b = (1/2) * l -------- (1)
2 * ((b + 5) + (l - 2)) = 60--------(2)
equation 2 gives l + b =27 -----(3)
substitute 1 in 3 and simplify to get the required values